Recent Developements in Inverse Problems

November 27 (Monday)- 30 (Thursday), 2000 Room 420 Research Institute of Mathematical Sciences

Program

November 27 (Monday) 14:20-18:00 14:20-15:20 M. I. Belishev (The Steklov Mathematical Institute, Russia) On reconstruction of metric via external electromagnetic measurements 15:40-16:40 Kazumi Tanuma (Osaka Kyoiku University) Direct determination of the derivatives of conductivity at the boundary from the localized Dirichlet to Neumann map 17:00-18:00 Takashi Suzuki (Osaka University) Exploitation and mathematical analysis of a new algorithm solver to inverse problems in magnetoencephalography

November 28 (Tuesday) 9:40-17:10 9:40-10:40 Choi Jong-Sung (University of Tokyo) Inverse source problems in the compressible viscous fluid dynamics 11:00-12:00 V. Isakov (Wichita University, USA) Inverse Problems for Quasilinear Equations 13:30-14:30 Jin Cheng (Gunma University) Inverse Problems of Diffractive Optics 14:50-15:50 Shigeru Sakaguchi (Ehime University) Stationary critical points of the heat flow and symmetries of the domains 16:10-17:10 Tan Yongji (Fudan University, People’s Republic of China) Parameter identification in resistivity well-loggings

November 29 (Wednesday) 9:40-17:45 9:40-10:40 Gen Nakamura (Gunma University) Reconstruction of obstacles with various type of boundary conditions for acoustic scattering 11:00-12:00 Masahiro Yamamoto (University of Tokyo) Global Lipschitz stability in an inverse hyperbolic problem by interior observation 13:30-14:30 I. Trooshin (Precision Mechanics and Control Institute, Russia) Inverse eigenvalue problem for a nonsymmetric system of ordinary differential equations of first order 14:50-15:50 Kiyoshi Mochizuki (Tokyo Metropolitan University) Spectral representations and scattering for w_{tt} + b(x)w_{t} − Δw = 0 16:10-17:10 Saburou Saitoh (Gunma University) Applications of analyticity of functions to inverse problems 17:15-17:45 MD. Asaduzzaman (Gunma University) Inversion formulas for the Reznitskaya transform and stability of Lipschitz type in determination of initial heat distribution

November 30 (Thursday) 9:40-11:30 9:40-10:10 Mishio Kawashita (Ibaraki University) Scattering theory by Lax and Phillips for the wave equation with infinite interface boundary 10:30-11:30 Masaru Ikehata (Gunma University) Obtaining information about the location of an unknown inclusion in the infinite strip

Reference: Masaru Ikehata Department of Mathematics, Faculty of Engineering, Gunma University

Tenjin-cho 1-5-1, Kiryu 376-8515, Japan (e-mail address) ikehata@math.sci.gunma-u.ac.jp